# Zero-Waiting Load Balancing with Heterogeneous Servers in Heavy Traffic

### v1

@@ Appendix D State Space Collapse in Super-Halfin-Whitt Regime
$=$ $\sum_{n=1}^{N}\lambda NA_{n}(q)(V(q+e_{n})-V(s))+\mu_{n}\mathbb{I}(q_{n}\geq 1)(V(q-e_{n})-V(q)).$
@@
$\Pr\left(V(q)\geq 3c_{1}N^{1-\alpha}\right)$ $\leq\left(\frac{N}{N+\frac{\epsilon k}{2(b-1)}N^{\alpha}\log N}\right)^{c_{1}N^{1-\alpha}}$
@@
$\Pr\left(V(q)\geq 3c_{1}N^{1-\alpha}\right)\leq\exp\left(-r\log N\right).$
@@ Appendix E Stein’s Method and Iterative Moment Bounds
$\mathbb{E}[Gf(Q)]=E\left[Gg\left(\bar{Q}\right)\right]=0.$ (33)
@@
$g(x+\frac{1}{N})-g\left(x\right)=$ $\left(g(x+\frac{1}{N})-g\left(x\right)\right)\left(1_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+1_{x>\eta+\frac{1}{N}}\right)$
$=$ $\frac{g^{\prime}(\xi)}{N}1_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+\left(\frac{g^{\prime}(x)}{N}+\frac{g^{\prime\prime}(\zeta)}{2N^{2}}\right)1_{x>\eta+\frac{1}{N}}$ (37)
$g(x-\frac{1}{N})-g\left(x\right)=$ $\left(g(x-\frac{1}{N})-g\left(x\right)\right)\left(1_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+1_{x>\eta+\frac{1}{N}}\right)$
$=$ $-\frac{g^{\prime}(\tilde{\xi})}{N}1_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+\left(-\frac{g^{\prime}(x)}{N}+\frac{g^{\prime\prime}(\tilde{\zeta})}{2N^{2}}\right)1_{x>\eta+\frac{1}{N}}$ (38)
@@ E.1 Gradient Bounds
where the last inequality holds because
$\left(\bar{Q}-\eta\right)^{r}\mathbb{I}_{\bar{Q}>\eta+\frac{1}{N}}\geq 0,$
Finally, we have

### v2

@@ Appendix D State Space Collapse in Super-Halfin-Whitt Regime
$=$ $\sum_{n=1}^{N}\lambda N\mathbb{P}_{n}(q)(V(q+e_{n})-V(q))+\mu_{n}\mathbb{I}(q_{n}\geq 1)(V(q-e_{n})-V(q)),$
where $\mathbb{P}_{n}(q)$ is the probability that an arriving job joins the server $n$ given the queue state $q.$
@@
$\mathbb{P}\left(V(Q)\geq 3c_{1}N^{1-\alpha}\right)$ $\leq\left(\frac{N}{N+\frac{\epsilon k}{2(b-1)}N^{\alpha}\log N}\right)^{c_{1}N^{1-\alpha}}$
@@
$\mathbb{P}\left(V(Q)\geq 3c_{1}N^{1-\alpha}\right)\leq\exp\left(-r\log N\right).$
@@ Appendix E Stein’s Method and Iterative Moment Bounds
$\mathbb{E}[Gf(Q)]=\mathbb{E}\left[Gg\left(\bar{Q}\right)\right]=0.$ (33)
@@
$g(x+\frac{1}{N})-g\left(x\right)=$ $\left(g(x+\frac{1}{N})-g\left(x\right)\right)\left(\mathbb{I}_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+\mathbb{I}_{x>\eta+\frac{1}{N}}\right)$
$=$ $\frac{g^{\prime}(\xi)}{N}\mathbb{I}_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+\left(\frac{g^{\prime}(x)}{N}+\frac{g^{\prime\prime}(\zeta)}{2N^{2}}\right)\mathbb{I}_{x>\eta+\frac{1}{N}}$ (37)
$g(x-\frac{1}{N})-g\left(x\right)=$ $\left(g(x-\frac{1}{N})-g\left(x\right)\right)\left(\mathbb{I}_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+\mathbb{I}_{x>\eta+\frac{1}{N}}\right)$
$=$ $-\frac{g^{\prime}(\tilde{\xi})}{N}\mathbb{I}_{\eta-\frac{1}{N}\leq x\leq\eta+\frac{1}{N}}+\left(-\frac{g^{\prime}(x)}{N}+\frac{g^{\prime\prime}(\tilde{\zeta})}{2N^{2}}\right)\mathbb{I}_{x>\eta+\frac{1}{N}}$ (38)
@@ E.1 Gradient Bounds
where the last inequality holds by dropping the negative term of $\lambda A_{b}(Q)$ and recalling the definition of $h(\bar{Q})=\max\{\bar{Q}-\eta,0\}$. Finally, we have